40m
QIL
Cryo_Lab
CTN
SUS_Lab
CAML
OMC_Lab
CRIME_Lab
FEA
ENG_Labs
OptContFac
Mariner
WBEEShop
|
PSL |
Not logged in |
 |
|
Tue Jul 2 11:21:24 2013, Erica, DailyProgress, fiber optic, mode matching 
|
Wed Jul 3 09:46:48 2013, Erica, DailyProgress, fiber optic, mode matching
|
Thu Jul 11 01:40:19 2013, Erica, DailyProgress, fiber optic, mode matching 
|
Fri Jul 12 00:51:39 2013, Erica, Notes, fiber optic, mode matching   
|
Tue Jul 16 10:27:47 2013, Erica, Notes, fiber optic, mode matching, beam recombination
|
Tue Jul 16 11:02:47 2013, Erica, Notes, fiber optic, mode matching, beam recombination
|
Wed Jul 17 11:58:47 2013, Erica, Notes, fiber optic, mode matching, beam recombination 
|
Wed Jul 17 18:29:37 2013, Erica, DailyProgress, fiber optic, reducing drift in the recombined beam
|
Sat Jul 20 23:39:09 2013, Erica, DailyProgress, fiber optic, reducing drift in the recombined beam
|
Tue Jul 23 10:29:12 2013, Erica, DailyProgress, fiber optic, reducing drift in the recombined beam 
|
Tue Jul 23 11:43:00 2013, Evan, DailyProgress, Electronics Equipment, Zach's op-amp recommendations
|
Fri Jul 26 11:07:44 2013, Erica, DailyProgress, BEAT, circuit for measuring temperature fluctations on CTN table   
|
Fri Jul 26 11:22:45 2013, Erica, DailyProgress, BEAT, circuit for measuring temperature fluctations on CTN table
|
Fri Jul 26 11:45:28 2013, Erica, DailyProgress, BEAT, circuit for measuring temperature fluctations on CTN table
|
Mon Jul 29 13:30:26 2013, Erica, DailyProgress, BEAT, adding resistor, capacitor, and sockets   
|
Tue Jul 30 12:23:41 2013, Erica, DailyProgress, BEAT, addressed noise problem, DC signal
|
Thu Aug 1 00:25:02 2013, Erica, DailyProgress, fiber optic, insulating foam  
|
Fri Aug 2 12:38:41 2013, tara, DailyProgress, fiber optic, insulating foam 
|
Tue Aug 6 12:31:24 2013, Erica, DailyProgress, fiber optic, insulating foam
|
Wed Aug 7 13:27:00 2013, Erica, DailyProgress, fiber optic, more foam, calculations for convertion from voltage to frequency  
|
Wed Aug 7 18:53:20 2013, Erica, DailyProgress, fiber optic, noise budget, Matlab notes. 
|
Tue Aug 13 12:39:36 2013, Erica, DailyProgress, fiber optic, rewrote calculations for converting from voltage noise to frequency noise  
|
Tue Aug 13 12:28:20 2013, Erica, Notes, Electronics Equipment, new AD 590 circuit and its noise calculation, beam alignment    
|
Fri Aug 16 10:35:44 2013, Erica, Notes, Electronics Equipment, AD590 msrmt, conversion to freq noise
|
Tue Jul 30 16:19:32 2013, Erica, Notes, fiber optic, setup for beam recombination 
|
Mon Aug 5 12:21:00 2013, Erica, Notes, fiber optic, calculation: phase change for fiber
|
Tue Jul 30 15:48:47 2013, Erica, DailyProgress, fiber optic, noise for GYRO
|
|
Message ID: 1251
Entry time: Tue Jul 23 10:29:12 2013
In reply to: 1245
Reply to this: 1252
1256
|
Author: |
Erica |
Type: |
DailyProgress |
Category: |
fiber optic |
Subject: |
reducing drift in the recombined beam |
|
|
July 22, 2013
We placed a plastic container from the GYRO lab over the spool of fiber. We wiped everything down w/ isopropanol. The cap is placed top side up and then the bottom is placed upside down on to the top of the cap. This is because the inside of the cap has ridges. There are two holes that the fiber is threaded thru. We used bubble wrap to cover it. It's not the most airtight but it should be okay. We left the aluminum foil surrounding the spool, to restrict air flow into the fiber. Then we left it there for awhile but coming back to it, there wasn't much improvement.
Evan suggested using a AD 590 to measure the temperature fluctuations of the table. Rana suggested we use the substitute for AD 743 that Zach found, which I need to find. Went to 40m to get a nicer circuit board and find a FET OpAmp with low current noise. We got a Vector Circboard 8015 and two OPA604's. We'll be building two setups so we can cross-correlate them and make sure the measurements are accurate.
tomorrow we will be soldering stuff.

 |